### v3

@@ 2.3 Jet injection
In our simulations, AGN jets are activated from the start. During the active phase (when $t\leq t_{\text{jet}}$), a constant jet is injected along the $+z$ direction, injecting mass, momentum, magnetic field, and energy into a cylindrical region with a cross-sectional radius of $R_{0}=1$ kpc and a height extending from $0\leq z\leq 1$ kpc along the $+z$ axis. The jet velocity is set at a constant value of $v_{\text{jet}}=0.1c$, where $c$ is the speed of light. The total energy injected into our simulated domain, representing the energy of one side of the jet, is $E_{\text{inj}}=0.5E_{\text{jet}}$, where $E_{\text{jet}}$ denotes the total energy of the bipolar jet system.
@@
where $t_{s}$ represents the typical sound crossing time scale in the ambient gas, which is calculated within a characteristic feedback radius $R_{\text{fb}}$. $R_{\text{fb}}$ is the radius within which the initial thermal energy of the ambient gas is equal to the jet outburst energy (Duan & Guo 2020). As $R_{\text{fb}}$  is determined by the initial gas distribution and jet injection energy, both of which are fixed in our simulations, it remains the same as 31 kpc for all our runs. In the context of this work, we have $t_{s}\approx 50\text{ Myr}$ and $P_{\text{fb}}\approx 1.3\times 10^{44}\text{ erg s}^{-1}$. The physical meaning of $P_{\text{fb}}$ lies in the fact that if the jet power, $P_{\text{jet}}$, significantly exceeds $P_{\text{fb}}$, an appreciable fraction of the power deposited by the jet will be dissipated in shocks. Furthermore, taking into account the Eddington luminosity of the black hole $L_{\text{Edd}}\approx 6.71\times 10^{46}\text{ erg s}^{-1}$, the ratio of jet power to Eddington luminosity is $\lambda_{\text{jet}}=P_{\text{jet}}/L_{\text{Edd}}$. For the case where $P_{\text{jet}}=P_{\text{fb}}$, this ratio yields $\lambda_{\text{jet}}\approx 2\times 10^{-3}$.
@@ 3.3 Radio power of the simulated sources
Figure 3: Evolution of the simulated radio sources in the P-D diagram. Distinct colors represent different powers, while different line styles signify distinct fractions of energy components and the magnetic structure parameter, $\alpha_{p}$, as indicated in the plot. The red shade covers regions with radio power one order of magnitude lower than that indicated by the red solid line. The detailed methodology for calculating radio power is described in Appendix C. The data points utilized in this analysis are sourced from Dabhade et al. (2020b).
@@ 4.1 Energy transport processes
We examine energy transport processes in different runs before discussing in detail the production of large radio sources. We now focus on two sets of runs: the runs with the same energy fractions ($f_{\rm m}=0.8$, $f_{\rm th}=0.05$, $\alpha_{\rm p}=0.3$) but different powers and the runs with the same power ($t_{\rm jet}=50\ \rm Myr$) but different energy fractions or magnetic field structure. The upper panels in Fig. 5 show the evolution of the distance that the jetted bubble traveled (half of the dimension of radio sources) in these two sets of runs. The evolution of the total energy residing in jetted bubbles is also shown in the lower panels of 5. As an approximation, the bubble region is identified by a density lower than $5\times 10^{-27}\rm{g\ cm^{-3}}$ and a toroidal magnetic field stronger than $10^{-9}\rm{G}$. The dotted vertical lines show when the same total energy is injected in different runs. We can see the scales of the jetted bubbles generally shows positive correlation with the total energy left in the bubbles particularly for the jets with the same energy makeup but different active times (powers). Notably, the high-power jet (run t5fm8) expels energy more rapidly upon jet cessation, leaving less energy within the bubble, which impedes the growth of the radio sources. Actually the most high-power jets can transmit energy outward through more potent shocks compared to low-power jets, which rely on less potent processes such as sound waves and slow expansions (Tang & Churazov 2017; Bambic & Reynolds 2019; Duan & Guo 2020).
@@
The evolution of these powers varying with the traveling distance of lobes is shown in Fig. 6. In all cases and in most regions, we can see that the energy transport processes are dominated by either thermal pressure or magnetic pressure. The thermal pressure power and magnetic pressure power in the higher power jet runs are both higher than those in lower power jet runs at the inner regions of clusters, as shown in the left panel. This means that higher power jets experience greater adiabatic losses in the inner regions of clusters, resulting in less energy residing in the bubbles, as discussed in the last paragraph. This should be related to the ”lower power-larger bubble” effect. On the other hand, we can see that lower power jets can transport more energy outside in the outer region of clusters, as indicated by the left panel of Fig. 6. Different energy makeups of jets do not result in as prominent differences as different powers, as shown in the right panel of Fig. 6.
@@ 4.2 Production of large radio sources
To gain some insight into the production of large radio sources, we examine the evolution of the widths of the jetted bubbles and show the results in the upper panels of Fig. 7. We can see that, within the central region of the halos (for example, r¡100kpc), the widths of the bubbles generally exhibit an anti-correlation with the ultimate extent of the simulated radio sources (see Figs. 1 and 5). Furthermore, we check the volume-averaged internal pressure (magnetic pressure plus thermal pressure) within the bubbles and present the results in the lower panels of Fig. 7. We can see that internal pressures of bubbles show a positive correlation with the widths of bubbles in the innermost region of the halos. The dotted vertical lines show where the correlations have just begun to develop. This phenomenon is due to the fact that higher internal pressures within bubbles will cause more prominent transverse expansion, resulting in larger cross sections. The larger cross sections will make the jets much less penetrating and easier to be decelerated by the dense gas in the center of the halos. In the upper left panel of Fig. 7, we can observe that higher power jets generally produce wider bubbles at the center of the halos. Subsequently, there will be stronger interactions between the higher power jets and the surrounding gas, causing them to transfer more energy than lower power jets in the center of the halos, as discussed above.
@@
On the other hand, the jetted bubbles with stronger poloidal magnetic field and thermal energy (run t50fm8bp3 and t50fth8) also seem to experience stronger Kelvin-Helmholtz instability than the bubbles jetted by toroidal magnetic field dominant jets, as shown in Fig. 10. This will make things more complicated, as the Kelvin-Helmholtz instability can trigger more prominent mixing with the surrounding medium. The mixing processes will actually make the low density bubbles lose more matter and energy than the stable ones, and finally reduce their potential to grow into large radio sources. This phenomenon is more pronounced in the case of higher power jets, as illustrated in Fig. 10. We will address more details regarding the relationship between the structure of the magnetic field and the Kelvin-Helmholtz instability in future studies.
Figure 6: Thermal pressure power (solid lines), magnetic pressure power (dashed lines) and magnetic tension power (dotted lines), calculated using Eqs. 10-12, varying with the traveling distance of lobes. Different colors represent the two sets of runs discussed in Sect. 4.1.
@@ Appendix A Gravitation and gas environment
Figure 8: Profiles of density and temperature are presented for the virial mass of $M_{\text{vir}}=10^{14}M_{\odot}$ (red solid lines, $f_{b}=0.1$, adopted in this work) and $M_{\text{vir}}=7.06\times 10^{14}M_{\odot}$ (blue solid lines, $f_{b}=0.14$), the latter being contrasted with the fitting results of galaxy cluster A1795 (blue dashed line, (Guo et al. 2018).
@@ Appendix B Injection of magnetic energy
We have developed a straightforward method to regulate the injection of magnetic energy of jets. Following Mbarek & Caprioli (2019), the toroidal magnetic field is set proportional to $r$ inside the jet base and $1/r$ outside, as given by:
$B_{\text{inj},\phi}=\left\{\begin{aligned} B_{0}\frac{r}{R_{0}},&\text{for }r<R_{0},\\ B_{0}\frac{R_{0}}{r},&\text{for }r>R_{0}.\end{aligned}\right.$ (26)
where $R_{0}$ is the cross-sectional radius of the jet base. Following Li et al. (2006) and Gan et al. (2017), the poloidal magnetic field components are defined as:
@@
where $\alpha_{p}$ is the parameter of the magnetic field structure that determines the ratio between the poloidal and toroidal magnetic fields. It is straightforward to prove that this magnetic field configuration satisfies $\nabla\cdot\textbf{B}=0$ analytically in cylindrical coordinates. The power of magnetic energy injection is defined as:
@@
Using Eqs. 26 - 30, we can analytically solve for $dB_{0}^{2}/dt$, followed by $dB_{\rm inj,r}^{2}/dt$, $dB_{\rm inj,z}^{2}/dt$, and $dB_{\rm inj,\phi}^{2}/dt$. Finally, the magnetic field is updated via the equation:
$\textbf{B}^{2}(t+\Delta t)=\textbf{B}^{2}(t)+\frac{d\textbf{B}^{2}}{dt}\Delta t.$ (31)
This method enables us to control the fraction of magnetic energy injected into the system, which is crucial for our work.
@@ Appendix C Calculation of synchrotron radiation
where $e$ represents the total energy density in Eq. 3. For self-consistency, we adopt $\eta_{e}=0.1$ as we expect the non-thermal electrons to have negligible impact on the dynamics of the jets and lobes in our simulations. Finally, $N_{0}$ can be derived from Eqs. 38 and 39. In Fig. 9, we show the volume-averaged energy density of various energy components varying with the traveling distance of the lobes.

### v4

@@ 2.3 Jet injection
In our simulations, AGN jets are activated from the start. During the active phase (when $t\leq t_{\text{jet}}$), a constant jet is injected along the $+z$ direction, injecting mass, momentum, and non-magnetic energy into a cylindrical region with a cross-sectional radius of $R_{0}=1$ kpc and extending from $z=0$ to a height of $z=h_{\text{\rm jet}}=1$ kpc along the $+z$ axis. The method for injecting magnetic field is described in Appendix B. The jet velocity is set at a constant value of $v_{\text{jet}}=0.1c$, where $c$ is the speed of light. The total energy injected into our simulated domain, representing the energy of one side of the jet, is $E_{\text{inj}}=0.5E_{\text{jet}}$, where $E_{\text{jet}}$ denotes the total energy of the bipolar jet system.
@@
where $t_{s}$ represents the typical sound crossing time scale in the ambient gas, which is calculated within a characteristic feedback radius $R_{\text{fb}}$. $R_{\text{fb}}$ is the radius within which the initial thermal energy of the ambient gas is equal to the jet outburst energy (Duan & Guo 2020). As $R_{\text{fb}}$  is determined by the initial gas distribution and jet injection energy, both of which are fixed in our simulations, it remains the same, at 31 kpc, for all our runs. In the context of this work, we have $t_{s}\approx 50\text{ Myr}$ and $P_{\text{fb}}\approx 1.3\times 10^{44}\text{ erg s}^{-1}$. The physical meaning of $P_{\text{fb}}$ lies in the fact that if the jet power, $P_{\text{jet}}$, significantly exceeds $P_{\text{fb}}$, an appreciable fraction of the power deposited by the jet will be dissipated in shocks. Furthermore, taking into account the Eddington luminosity of the black hole $L_{\text{Edd}}\approx 6.71\times 10^{46}\text{ erg s}^{-1}$, the ratio of jet power to Eddington luminosity is $\lambda_{\text{jet}}=P_{\text{jet}}/L_{\text{Edd}}$. For the case where $P_{\text{jet}}=P_{\text{fb}}$, this ratio yields $\lambda_{\text{jet}}\approx 2\times 10^{-3}$.
@@ 3.3 Radio power of the simulated sources
Figure 3: Evolution of the simulated radio sources in the P-D diagram. Distinct colors represent different powers, while different line styles signify distinct fractions of energy components and the magnetic structure parameter, $\alpha_{p}$, as indicated in the plot. The blue shade covers regions with radio power one order of magnitude lower than that indicated by the blue solid line. The detailed methodology for calculating radio power is described in Appendix C. The data points utilized in this analysis are sourced from Dabhade et al. (2020b).
@@ 4.1 Energy transport processes
We examine energy transport processes in different runs before discussing in detail the production of large radio sources. We now focus on two sets of runs: the runs with the same energy fractions ($f_{\rm m}=0.8$, $f_{\rm th}=0.05$, $\alpha_{\rm p}=0.3$) but different powers and the runs with the same power ($t_{\rm jet}=50\ \rm Myr$) but different energy fractions or magnetic field structure. The upper panels in Fig. 5 show the evolution of the distance that the jetted bubble traveled (half of the dimension of radio sources) in these two sets of runs. The evolution of the total energy residing in jetted bubbles is also shown in the lower panels of 5. As an approximation, the bubble region is identified by a density lower than $5\times 10^{-27}\rm{g\ cm^{-3}}$ and a toroidal magnetic field stronger than $10^{-9}\rm{G}$. The dotted vertical lines show when the fixed total energy has been injected and the jet is turned off in different runs. We can see the scales of the jetted bubbles generally shows positive correlation with the total energy left in the bubbles particularly for the jets with the same energy makeup but different active times (powers). Notably, the high-power jet (run t5fm8) expels energy more rapidly upon jet cessation, leaving less energy within the bubble, which impedes the growth of the radio sources. Actually the most high-power jets can transmit energy outward through more potent shocks compared to low-power jets, which rely on less potent processes such as sound waves and slow expansions (Tang & Churazov 2017; Bambic & Reynolds 2019; Duan & Guo 2020).
@@
The Maxwell stress power can be expressed as $P_{out,m}=P_{out,m1}+P_{out,m2}$. The evolution of thermal pressure power, Maxwell stress power and magnetic pressure power varying with the traveling distance of lobes is shown in Fig. 6. The magnetic tension power can be determined by the difference between the Maxwell stress power and the magnetic pressure power, which can be negative for some cases (for example, during the early stage of run t50fm8bp3, as shown in the right panel of Fig. 6). The thermal pressure power and Maxwell stress power in the higher power jet runs are both higher than those in lower power jet runs at the inner regions of clusters, as shown in the left panel. This means that higher power jets experience greater adiabatic losses in the inner regions of clusters, leaving less energy residing in the bubbles at later times, as discussed in the last paragraph. This should be related to the ”lower power-larger bubble” effect. On the other hand, we can see that lower power jets can transport more energy outside in the outer region of clusters, as indicated by the left panel of Fig. 6. Different energy makeups of jets do not result in as prominent differences as different powers, as shown in the right panel of Fig. 6.
@@ 4.2 Production of large radio sources
To gain some insight into the production of large radio sources, we examine the evolution of the widths of the jetted bubbles and show the results in the upper panels of Fig. 7. We can see that, within the central region of the halos (for example, r¡100kpc), the widths of the bubbles generally exhibit an anti-correlation with the ultimate extent of the simulated radio sources (see Figs. 1 and 5). Furthermore, we check the volume-averaged internal pressure (magnetic pressure plus thermal pressure) within the bubbles and present the results in the lower panels of Fig. 7. We can see that internal pressures of bubbles show a positive correlation with the widths of bubbles in the innermost region of the halos. The dotted vertical lines show where the correlations have just begun to develop. This phenomenon is due to the fact that higher internal pressures within bubbles will cause faster transverse expansion, resulting in larger cross sections. The larger cross sections will make the jets much less penetrating and easier to be decelerated by the dense gas in the center of the halos. In the upper left panel of Fig. 7, we can observe that higher power jets generally produce wider bubbles at the center of the halos. Subsequently, there will be stronger interactions between the higher power jets and the surrounding gas, causing them to transfer more energy than lower power jets in the center of the halos, as discussed above.
@@
On the other hand, the inflated bubbles with stronger poloidal magnetic field and thermal energy (run t50fm8bp3 and t50fth8) also seem to experience stronger Kelvin-Helmholtz instability than the bubbles jetted by toroidal magnetic field dominant jets, as shown in Fig. 10. This will make things more complicated, as the Kelvin-Helmholtz instability can trigger more prominent mixing with the surrounding medium. The mixing processes will actually make the low density bubbles lose more matter and energy than the stable ones, and finally reduce their potential to grow into large radio sources. This phenomenon is more pronounced in the case of higher power jets, as illustrated in Fig. 10. We will address more details regarding the relationship between the structure of the magnetic field and the Kelvin-Helmholtz instability in future studies.
Figure 6: Thermal pressure power (solid lines), Maxwell stress power (dashed lines) and magnetic pressure power (dotted lines), calculated using Eqs. 10-12, varying with the traveling distance of lobes. Different colors represent the two sets of runs discussed in Sect. 4.1.
@@ Appendix A Gravitation and gas environment
Figure 8: Profiles of density and temperature are presented for the virial mass of $M_{\text{vir}}=10^{14}M_{\odot}$ (orange solid lines, $f_{b}=0.1$, adopted in this work) and $M_{\text{vir}}=7.06\times 10^{14}M_{\odot}$ (blue solid lines, $f_{b}=0.14$), the latter being contrasted with the fitting results of galaxy cluster A1795 (blue dashed line, (Guo et al. 2018).
@@ Appendix B Injection of magnetic energy
We have developed a straightforward method to regulate the injection of magnetic energy of jets. Following Mbarek & Caprioli (2019) with modifications, the toroidal magnetic field is set proportional to $r$ inside the jet base, $1/r$ outside and zero further away, as given by:
$B_{\text{inj},\ \phi}=\left\{\begin{aligned} B_{0}\frac{r}{R_{0}},\ \ &\text{for }r<R_{0}\ \text{and}\ z<h_{\rm jet},\\ B_{0}\frac{R_{0}}{r},\ \ &\text{for }R_{0}<r<2R_{0}\ \text{and}\ z<h_{\rm jet},\\ 0,\ \ &\text{for }r>2R_{0}\ \text{or}\ z>h_{\rm jet},\\ \end{aligned}\right.$ (26)
where $R_{0}$ and $h_{\rm jet}$ are the cross-sectional radius and height of the jet base respectively. Following Li et al. (2006) and Gan et al. (2017), the poloidal magnetic field components are defined as:
@@
where $\alpha_{p}$ is the parameter of the magnetic field structure that determines the ratio between the poloidal and toroidal magnetic fields. The poloidal magnetic field is injected throughout the entire simulation region, although its value is negligible for cells far from the jet base. It is straightforward to prove that this magnetic field configuration satisfies $\nabla\cdot\textbf{B}=0$ analytically in cylindrical coordinates. Then, the injection rate of magnetic energy can be expressed as:
$\dot{e}_{\text{b},\ \phi}=\left\{\begin{aligned} \frac{1}{2}\dot{e}_{b0}\frac{r^{2}}{R_{0}^{2}},\ \ &\text{for }r<R_{0}\ \text{and}\ z<h_{\rm jet},\\ \frac{1}{2}\dot{e}_{b0}\frac{R_{0}^{2}}{r^{2}},\ \ &\text{for }R_{0}<r<2R_{0}\ \text{and}\ z<h_{\rm jet},\\ 0,\ \ &\text{for }r>2R_{0}\ \text{or}\ z>h_{\rm jet},\\ \end{aligned}\right.$ (29)
$\dot{e}_{\text{b},\ z}=\frac{1}{2}\alpha_{p}^{2}\dot{e}_{b0}\cdot 4\left(1-\frac{r^{2}}{R_{0}^{2}}\right)^{2}\exp\left(-2\frac{r^{2}+z^{2}}{R_{0}^{2}}\right),$ (30)
$\dot{e}_{\text{b},\ r}=\frac{1}{2}\alpha_{p}^{2}\dot{e}_{b0}\cdot\frac{4z^{2}r^{2}}{R_{0}^{4}}\exp\left(-2\frac{r^{2}+z^{2}}{R_{0}^{2}}\right),$ (31)
where $\dot{e}_{b0}=\frac{d}{dt}(B_{0}^{2})$. The power of magnetic energy injection is defined as:
@@
$=[\ (\frac{1}{4}+\text{ln}2)\pi h_{\rm jet}R^{2}_{0}+\frac{5\pi}{16}\sqrt{\frac{\pi}{2}}\alpha^{2}_{p}R^{3}_{0}]\ \dot{e}_{b0}$
@@
Combining Eqs. 32 and 33 the value of $\dot{e}_{b0}$ can be determined. Then, using Eqs. 29 - 31, the injection rate of magnetic energy can be solved for all regions. The magnitude of magnetic field is updated via the equation:
$\frac{\partial B_{i}^{2}}{\partial t}=2\dot{e}_{\text{b},\ i},$ (34)
where $i$ represents $\phi$, $z$ or $r$. This means that the difference form of the injection rate of the magnetic field can be expressed as:
$\frac{B_{i}(t+\Delta t)-B_{i}(t)}{\Delta t}\approx\frac{\pm[B^{2}_{i}(t)+2\dot{e}_{\text{b},\ i}\Delta t]^{\frac{1}{2}}-B_{i}(t)}{\Delta t}.$ (35)
The sign $\pm$ is controlled to be consistent with Eq.26-28, which means $B_{z}$ is injected positively at $r<R_{0}$ and negatively at $r>R_{0}$, whereas $B_{\phi}$ and $B_{r}$ are both injected positively throughout the entire injection region. This method enables us to control the fraction of magnetic energy injected into the system, which is crucial for our work. To be consistent with Eq.4, Eq.34 would require:
$\dot{B}_{\text{inj},\ i}=\frac{\dot{e}_{\text{b},\ i}}{B_{i}}.$ (36)
@@ Appendix C Calculation of synchrotron radiation
where $e$ represents the total energy density in Eq. 3. For self-consistency, we adopt $\eta_{e}=0.1$ as we expect the non-thermal electrons to have negligible impact on the dynamics of the jets and lobes in our simulations. Finally, $N_{0}$ can be derived from Eqs. 43 and 44. In Fig. 9, we show the volume-averaged energy density of various energy components varying with the traveling distance of the lobes.